Tickets for a play at the community theater cost $20 for an adult and $4 for a child. If 140 tickets were sold and the total receipts were $2160, how many of each type of ticket were sold?

Answer by elima(1433)

Tickets for a play at the community theater cost $20 for an adult and $4 for a child. If 140 tickets were sold and the total receipts were $2160, how many of each type of ticket were sold?
------------------------------------------------------------
x=adult tickets
y=child tickets
x+y=140
20x+4y=2160
----------------
x=140-y
20(140-y)+4y=2160
2800-20y+4y=2160
-16y=-640
y=40
------------
x+40=140
x=140-40
x=100
------------------
so the were 100 adult tickets and 40 child tickets sold
:)